OpenStudy (anonymous):

Use a system of equations to solve the following problem: Charles High School has a total of 94 boys and girls who play sports. If the number of boys is 22 more than twice the number of girls, how many boys play sports at this high school? 70 48 24 80

OpenStudy (goformit100):

24

OpenStudy (anonymous):

no 70

OpenStudy (goformit100):

no war here , tell how

OpenStudy (anonymous):

et x = number of girls therefore, 22 + 2x = number of boys x + (22 + 2x) = 94 22 + 3x = 94 3x = 72 x = 24 girls 22 + 2x = 22 + 2(24) = 22 + 48 = 70 boys

OpenStudy (goformit100):

0k..... Thanks for guiding me..

OpenStudy (mayankdevnani):

answer is 24 @rebecca1233

OpenStudy (goformit100):

@ParthKohli who is correct ?

OpenStudy (mayankdevnani):

24 is girls

OpenStudy (goformit100):

I think i was correct